OpenStudy (anonymous):

write an equation for the line that passes through the point (-2,-3) with a slope of 1/2

OpenStudy (mathmale):

This strongly suggests that you use the point-slope form of the equation of a straight line. Please type that equation out in its general form, and then we'll apply it to this situation.

OpenStudy (anonymous):

y=mx+b

OpenStudy (mathmale):

No, that's the slope-intercept form. Please do a quick Internet search for : point-slope form definition

OpenStudy (anonymous):

The equation of a line is \[y=mx+b\] Plugging in the given info gives us \[-3=\frac{1}{2}(-2)+b\] Solving for b gives: \[b=-2\] Put everything into the equation to give \[y=\frac{1}{2}x-2\] Check: \[y=\frac{1}{2}x-2\] \[-3=\frac{1}{2}(-2)-2\] \[-3=-3\] CHECK